OpenStudy (anonymous):

I have filled a bag with differently colored ping-pong balls. There are twenty balls in the bag. Twelve are white, six are orange and two are pink. When I draw a ball from the bag, I remove it. What is the probability that my first draw is a white ball and my second is not? A) 0.240 B) 0.253 C) 0.400 D) 0.421

OpenStudy (raden):

it means asking to P(White and Orange) + P(White and pink) = 12/20 * 6/19 + 12/20 * 2/19 = 72/380 + 24/380 = 96/380 = 0.2526.. = 0.253

OpenStudy (anonymous):

Thanks so much

OpenStudy (raden):

you're welcome :)